# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v4

@@ IV-C Empirical Study Design
All configurations share the same injection model, prompt formatting, and RAG setup. We evaluate these variants on the same benchmark split (see Section 3, A) and report both formal and manual success metrics (see Section 3, D). This incremental comparison enables a fine-grained understanding of how individual agents affect the correctness, realism, and efficiency of the injection process.
@@ IV-D Evaluation and Metrics
We assess the success rate of vulnerability injection, as it directly indicates the expected label accuracy of datasets produced by our framework
@@ V Results and Analysis
Table III compares AVIATOR with other techniques, showing that its success rate on FormAI (91%) and PrimeVul (94%) surpasses existing injection methods, like VinJ and VulGen, which report 69% success rates. AVIATOR’s guided injection process allows for precise, context-aware edits that mimic developer-like mistakes, making it effective even in partially defined or real-world settings.
@@
It is important to note that, even without fine-tuning, AVIATOR achieves strong baseline performance. This indicates that the agentic workflow itself, through decomposition, correction loops, and validation, is sufficient to generate realistic and diverse vulnerabilities. Fine-tuning provides further gains, but is not strictly necessary for the framework to operate effectively. While baseline performance depends on how well the underlying LLM has been trained on the target language, this property remains especially valuable when extending to programming languages or vulnerability classes where aligned training data is scarce, since AVIATOR can still produce rich vulnerability patterns.
Learning dynamics
Refer to caption Refer to caption
[…]
Answer to RQ2. Fine-tuning the vulnerability injection agent using SFT leads to consistent and significant performance gains for the whole AVIATOR workflow, particularly on complex datasets like FormAI. GRPO fails to match this improvement despite being substantially more resource-intensive. These results highlight the efficiency and generalization strength of SFT, even when trained on an entirely separate dataset. Additionally, the results show strong baseline performance without fine-tuning, indicating that the agentic workflow alone can inject vulnerabilities. This is essential for transfer to new programming languages or CWE categories where aligned data is unavailable.
@@
However, figure 4 also reveals performance variability primarily on vulnerability categories that either lack precise injection semantics or require constraint satisfaction beyond the function scope. For instance, AVIATOR fails to inject CWE-391 (Unchecked Error Condition), which yields a 0% success rate. This CWE lacks a precise modern definition and is slated for deprecation, making it difficult to formulate a concrete injection objective. Similarly, CWE-190 (Integer Overflow) poses a greater challenge, achieving a 67% success rate, because successful injection often depends on satisfying strict arithmetic constraints that involve global assumptions or external inputs. These cases reflect inherent limitations of function-scoped injection rather than shortcomings of AVIATOR’s workflow.
@@ V-E Compute and Cost Estimates
To assess AVIATOR’s practical utility, we evaluate how augmenting training data with AVIATOR-generated vulnerable samples impacts downstream deep learning-based vulnerability detection (DLVD) performance. Table VII reports the performance metrics for four distinct deep learning architectures (Devign, Reveal, LineVul, CodeT5) trained on the Devign and PrimeVul datasets augmented with generated samples using different strategies.
[…]
TABLE VI: Overall Average Relative Improvement of AVIATOR Compared to Other Strategies.
Compared Strategy Precision Recall F1-Score
No Aug +12.2% +104.9% +21.8%
ROS +9.0% +26.0% +12.5%
VulGen +30.3% +5.7% +25.2%
VGX +39.2% -25.4% +24.6%
VulScribeR +1.4% +7.4% +2.9%
TABLE VII: Downstream vulnerability detection performance (Precision, Recall, F1). Each model is trained on the seed dataset augmented with 5k injected vulnerable samples using each strategy. Best F1 scores per model and setting are highlighted in bold. Results for non-AVIATOR strategies are taken from the VulScribeR paper [10]; we report new results for AVIATOR and CodeT5.
@@
When compared to pattern-based injection methods, AVIATOR improves average F1 score by +25.2% over VulGen and by +24.6% over VGX. Among all baselines, VGX is the only method that achieves a higher average recall than AVIATOR; however, this comes at the cost of a substantial drop in precision. This behavior is consistent with VGX’s reliance on single-statement injection templates and its limited coverage of vulnerability classes (23 CWEs). Such augmentation strategies can cause detectors to become overly sensitive (high recall) while substantially increasing false positives (low precision) which ultimately limits their practical utility for large-scale data augmentation [10]. In contrast, AVIATOR improves F1 through a more balanced precision-recall profile, indicating that it provides a cleaner and less artifact-driven training signal.
@@
Distributional fidelity and realism
AVIATOR achieves the lowest KL divergence under both metrics, which confirms that it remains closest to the original distribution. VulGen and VGX introduce notably larger shifts, consistent with augmentation that concentrates on a narrower set of injection patterns. VulScribeR reduces this distortion compared to pattern-mining baselines but still deviates more than AVIATOR. Overall, these distributional results align with the precision-recall trends in Table VI: lower distortion corresponds to more realistic, learnable injected vulnerabilities that improve generalization without inducing excessive false alarms.
@@ VII CONCLUSION
Data will be made available on request.

### v5

# AVIATOR: Towards AI-Agentic Vulnerability Injection Workflow for High-Fidelity, Large-Scale Code Security Dataset

@@ IV-C Empirical Study Design
All configurations share the same injection model, prompt formatting, and RAG setup. We evaluate these variants on the same benchmark split (see Section IV-A) and report both formal and manual success metrics (see Section IV-D). This incremental comparison enables a fine-grained understanding of how individual agents affect the correctness, realism, and efficiency of the injection process.
@@ IV-D Evaluation and Metrics
We assess the success rate of vulnerability injection, as it directly indicates the expected label accuracy of datasets produced by our framework.
@@ V Results and Analysis
It is important to note that, even without SFT, AVIATOR achieves strong baseline performance. This indicates that the agentic workflow, through decomposition, correction loops, and validation, is sufficient to generate realistic and diverse vulnerabilities. Fine-tuning provides further gains but is not required for the framework to operate effectively. While baseline performance depends on how well the underlying LLM has been trained on the target language, this property remains especially valuable when extending to programming languages or vulnerability classes where aligned training data is scarce, since AVIATOR can still produce rich vulnerability patterns.
Refer to caption Refer to caption
[…]
Learning dynamics
Answer to RQ2. Fine-tuning the vulnerability injection agent using SFT leads to consistent and significant performance gains for the whole AVIATOR workflow, particularly on complex datasets like FormAI. GRPO fails to match this improvement despite being substantially more resource intensive. These results highlight the efficiency and generalization strength of SFT, even when trained on an entirely separate dataset. Additionally, the results show strong baseline performance without fine-tuning, indicating that the agentic workflow alone can inject vulnerabilities. This is essential for transfer to new programming languages or CWE categories where aligned data is unavailable.
@@
However, figure 4 also reveals performance variability primarily on vulnerability categories that either lack precise injection semantics or require constraint satisfaction beyond the function scope. For instance, AVIATOR fails to inject CWE-391 (Unchecked Error Condition), which yields a 0% success rate. This CWE lacks a precise modern definition and is slated for deprecation, making it difficult to formulate a concrete injection objective. Similarly, CWE-190 (Integer Overflow) poses a greater challenge, achieving a 67% success rate, because successful injections often depend on satisfying strict arithmetic constraints that involve global assumptions or external inputs. These cases reflect inherent limitations of function-scoped injections rather than shortcomings of AVIATOR’s workflow.
@@ V-E Compute and Cost Estimates
To assess AVIATOR’s practical utility, we evaluate how augmenting training data with AVIATOR-generated vulnerable samples impacts downstream deep learning-based vulnerability detection (DLVD) performance. Table VI reports the performance metrics for four distinct deep learning architectures (Devign, Reveal, LineVul, CodeT5) trained on the Devign and PrimeVul datasets augmented with generated samples using different strategies.
TABLE VI: Downstream vulnerability detection performance (Precision, Recall, F1). Each model is trained on the seed dataset augmented with 5k injected vulnerable samples using each strategy. Best F1 scores per model and setting are highlighted in bold. Results for non-AVIATOR strategies are taken from the VulScribeR paper [10]; we report new results for AVIATOR and CodeT5.
@@
TABLE VII: Overall Average Relative Improvement of AVIATOR Compared to Other Strategies.
Compared Strategy Precision Recall F1-Score
No Aug +12.2% +104.9% +21.8%
ROS +9.0% +26.0% +12.5%
VulGen +30.3% +5.7% +25.2%
VGX +39.2% -25.4% +24.6%
VulScribeR +1.4% +7.4% +2.9%
[…]
@@
When compared to pattern-based injection methods, AVIATOR improves average F1 score by +25.2% over VulGen and by +24.6% over VGX. Among all baselines, VGX is the only method that achieves a higher average recall than AVIATOR; however, this comes at the cost of substantially lower precision. This behavior aligns with VGX’s reliance on single-statement injection templates and its limited coverage of vulnerability classes (23 CWEs). Such augmentation strategies can make detectors overly sensitive, yielding high recall but low precision (many false positives) which ultimately limits their practical utility for large-scale data augmentation [10]. In contrast, AVIATOR improves F1 through a more balanced precision-recall profile, indicating that it provides a cleaner and less artifact-driven training signal.
@@
Distributional fidelity and realism
AVIATOR achieves the lowest KL divergence under both metrics, which confirms that it remains closest to the original distribution. VulGen and VGX introduce notably larger shifts, consistent with augmentation that concentrates on a narrower set of injection patterns. VulScribeR reduces this distortion compared to pattern-mining baselines but still deviates more than AVIATOR. Overall, these distributional results align with the precision-recall trends in Table VII: lower distortion corresponds to more realistic, learnable injected vulnerabilities that improve generalization without inducing excessive false alarms.
@@ VII CONCLUSION
Data will be made available upon request.